Of 2000 persons under investigation (PUI) due


to Covid-19, 1050 PUI undergone rapid


testing, 500 PUI undergone swabtesting, and


200 PUI undergone both rapid and swab


testing. What is the probability that a


randomly selected PUI undergone random


testing or swab testing?


1
Expert's answer
2022-02-24T17:54:18-0500

Let "R" denote rapid testing, "S" denote swab testing.

Given

"P(R)=\dfrac{1050}{2000}=0.525, P(S)=\dfrac{500}{2000}=0.25,"

"P(R\cap S)=\dfrac{200}{2000}=0.1"

The probability that a randomly selected PUI undergone random testing or swab testing is

"P(R\cup S)=P(R)+P(S)-P(R\cap S)"

"=0.525+0.25-0.1=0.675"